THE EMISSION SPECTRA OF THE Ã STATE OF THE C_{3}-Xe VAN DER WAALS COMPLEX

JUN-MEI CHAO, KENG SENG THAM, GUIQIU ZHANG, ANTHONY J. MERER, and \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 10617, R. O. C..

Emission bands of the C_{3}-Xe van der Waals complex near the 2_{0}^{2−}, 2_{0}^{2+}, and 2_{0}^{4−} bands of the Ã−X̃ system of C_{3} have been recorded, but as yet definitive spectral assignments could not be made. Unlike those of the C_{3}-Ar and C_{3}-Kr complexes, the observed C_{3}-bending vibrational progressions of the Xe-complex could not be fitted to a simple dipole-induced dipole model potential^{a}. Ab initio calculations of the Ar- and Xe-complexes have been carried out using the CCSD(T) method at the cc-pVQZ level. Tentative vibrational assignments of the Xe-complex based upon the ab initio calculation will be discussed.

aG. Zhang, B.-G. Lin, S.-M. Wen, and Y.-C. Hsu, J. Chem. Phys. \underline{120}, 3189, 2004.